OpenStudy (anonymous):

The sum of three consecutive natural numbers is 1122, find the numbers. The three numbers in increasing order are

OpenStudy (yoloshroom):

x+1+x+2+x+3 i believe would work. so 3x +6 = 1122 3x=1116 x=372 so 373,374,375

OpenStudy (dn_20):

We have x + (x+1) + (x+2) = 1122 3x +3 = 1122 3x = 1119 x = 373 the three munber is 373, 374, 375 there is another way to solve this by finding the avg and then -1 for the first one and +1 for the third one